OpenStudy (anonymous):

A and B do work in 15 days B and C do it in 20 days if B can do it in 60 days how long can A and C do it together?

OpenStudy (anonymous):

(A + B)'s 1 day work = \(\frac{1}{15}\) (B + C)'s 1 day work = \(\frac{1}{20}\) B alone 1 day work = \(\frac{1}{60}\) Add all: (A + 2B + C)'s 1 day work = \(\frac{1}{15} + \frac{1}{20} + \frac{1}{60}\) Now subtract B's 1 day work: (A + 2B + C - 2B)'s 1 day work = \(\frac{1}{15} + \frac{1}{20} + \frac{1}{60} - \frac{2}{60}\) Solving it we get: (A + C )'s 1 day work = \(\frac{1}{10}\) So, (A + C) both will finish in 10 days..

OpenStudy (anonymous):

oh great thanx that makes sense. :)

OpenStudy (anonymous):

(A + 2B + C)'s 1 day work = 1/15+1/20 NOT '1/15+1/20+1/60'

OpenStudy (anonymous):

wait there's choices A) 15 days B) 17 days C) 16 days D) 18 days E) 12 days

OpenStudy (anonymous):

Wait I have done it wrong.. There is one more method..

OpenStudy (anonymous):

yeah i think u need to take it slow

OpenStudy (anonymous):

The answer is 'E'

OpenStudy (anonymous):

so after adding A and C wat dyu do nxt to get 12 days?

OpenStudy (anonymous):

(A + B)'s 1 day work = \(\frac{1}{15}\) So subtract B's 1 day work form it: A's 1 day work = \[\frac{1}{15} - \frac{1}{60} = \frac{3}{60}\] Similarly, (B + C)'s 1 day work = \(\frac{1}{20}\) Subtract B's 1 day work here: C's 1 day work = \[\frac{1}{20} - \frac{1}{60} = \frac{2}{60}\] Now add both: (A + C)'s 1 day work: \(\frac{3}{60} + \frac{2}{60} = \frac{5}{60} = frac{1}{12}\) (A + C)'s will complete in 12 days..

OpenStudy (anonymous):

(A + C)'s 1 day work: \[\frac{3}{60} + \frac{2}{60} = \frac{5}{60} = \frac{1}{12}\] So, (A+C)'s will complete it in 12 days..

OpenStudy (anonymous):

just like what waterineyes wrote

OpenStudy (anonymous):

right i think im startin to get it now